If angle RMK=5x+8, angle KMT=3x-5, and angle TML=6x+17, then x=_______
a. 10
b. 4
c. 5
d. 12

Answer by checkley71(8403)

IF THESE ARE COMPLEMENTARY ANGLES THEN THEIR SUM=90 DEGREES.
5X+8+3X-5+6X+17=90
14X+20=90
14X=70
X=70/14
X=5 DEGREES ANSWER.
IF THEY ARE SUPPLIMENTARY ANGLES THEIR SUM=180 DEGREES.
5X+8=3X-5+6X+17=180
14X+20=180
14X=180-20
X=160/14
X=11.42857 DEGREES.